-What is the simple interest rate on a $1050 investment paying $738.15 interest in 19 years?

-Chasity wants $9,000 saved in 3 years to make a down payment on a house. How much money should she invest now at 5.2% compounded annually in order to meet her goal?
Answer #1

A B 1050 738.15 Investment Interest Years Annual Interest Annual rate 19 38.85 3.70% 9000 Future value Years Rate Money to be

А в ср Investment Interest Years Annual Interest Annual rate 1050 738.15 19 =D6/07 =D8/D5 Future value 9000 Years Rate 0.052